# Rethinking Irregular Time Series Forecasting: A Simple yet Effective Baseline

### v1

@@ 3.4 Query-based Weighted Aggregation
After TAPA generates regularized patch representation sequences $H^{n}=[h^{n}_{1},\dots,h^{n}_{P}]\in\mathbb{R}^{P\times D}$ for each variable, APN employs a concise mechanism to summarize them into a single representation $H_{c}^{n}$, obviating complex inter-sequence transformation modules. To preserve the sequential order of patches, standard positional encodings $PE\in\mathbb{R}^{P\times D}$ are added to $H^{n}$, yielding position-aware patch representations $H_{pe}^{n}=H^{n}+PE$. Then, to dynamically select and fuse information most relevant to the prediction task, a learnable query vector $q^{n}\in\mathbb{R}^{D}$ is introduced for each variable $n$. This query vector interacts with all $P$ position-aware patches $H_{pe}^{n}$, calculating unregularized importance scores $s_{p}^{n}$ via dot product, which are then regularized into weight coefficients $\beta^{n}=[\beta_{1}^{n},\dots,\beta_{P}^{n}]$ using the Softmax function. Finally, the representation $H_{c}^{n}$ of variable $n$ is obtained by a weighted sum of all patches $h_{pe,p}^{n}$:
@@
where the summary representations of all variables $[H_{c}^{1},\dots,H_{c}^{N}]$ form the final representation matrix $H_{final}\in\mathbb{R}^{N\times D}$, which is then processed by a layer normalization to stabilize subsequent MLP input. This learnable Query-based Weighted Aggregation is concise and efficient, avoiding complex key-value projections or pairwise inter-patch interactions, demonstrating APN’s simple yet effective design.
@@ 3.5 Forecasting Decoder
After obtaining the summary historical representation $H_{final}^{n}\in\mathbb{R}^{D}$ for each variable $n$ and the target query time point $\tau_{k}$, APN employs a two-layer MLP decoder for prediction. The decoder takes the concatenation of the variable’s summary representation $H_{final}^{n}$ and the learnable temporal encoding $TE(\tau_{k})$ of the query time point $\tau_{k}$ as input, outputting the predicted value $\hat{v}^{n}_{k}$:
@@
3.6 Loss Function
The overall training objective of the APN model is to minimize the Mean Squared Error (MSE) between predicted and true observed values. For given historical observations $\mathcal{O}_{hist}$ and a set of query time points $\mathcal{T}_{pred}=\{\tau_{k}\}_{k=1}^{H}$, the loss function $\mathcal{L}$ is defined as:
$\mathcal{L}=\frac{1}{N\cdot H}\sum_{n=1}^{N}\sum_{k=1}^{H}(\hat{v}^{n}_{k}-v^{n}_{k})^{2},$ (10)
where $v^{n}_{k}$ is the true value of variable $n$ at query time $\tau_{k}$, and $\hat{v}^{n}_{k}$ is the model’s prediction.
@@ 4.1 Setup
Implementation Details: We implement APN and all related experiments using the PyTorch framework on a server equipped with an Intel(R) Xeon(R) Gold 6326 CPU (@ 2.90GHz) and an NVIDIA GeForce RTX 3090 GPU. For our APN model, we set the hidden dimension $D$ according to different datasets. Specifically, the hidden dimension $D$ is set to 64 for the PhysioNet and USHCN, while it is set to 16 for the Human Activity and MIMIC. The time encoding dimension $D_{te}$ is uniformly set to 10. The number of patches $P$ is adjusted based on the characteristics of the datasets: 20 for PhysioNet, 16 for MIMIC, 75 for Human Activity, and 4 for USHCN. The training batch size is set to 256 for all datasets. The model is trained using the Adam optimizer with an initial learning rate of $1\times 10^{-2}$. We train the model for a maximum of 200 epochs and employ an early stopping strategy with a patience of 50. All experiments are repeated 5 times, and the mean and standard deviation of the performance metrics are reported to ensure the robustness of the results.
@@
Metric MSE×$\times 10^{-3}$ MAE×$\times 10^{-2}$ MSE×$\times 10^{-2}$ MAE×$\times 10^{-2}$ MSE×$\times 10^{-3}$ MAE×$\times 10^{-2}$ MSE×$\times 10^{-1}$ MAE×$\times 10^{-1}$
@@ 4.2 Main results
We compare the APN model with nineteen baseline models on four challenging datasets—see Table 1. We have the following observations: 1) APN achieves leading prediction accuracy across the board. On all datasets, APN achieves optimal or highly competitive forecasting performance. Compared to the second-best performing model, T-PATCHGNN, APN achieves significant reductions in MSE and MAE metrics by approximately 9.63% and 5.24%, respectively. 2) APN demonstrates exceptional cross-domain generalization capability and robustness. Whether on IMTS datasets with different characteristics in healthcare (PhysioNet, MIMIC), biomechanics (Human Activity), or climate science (USHCN), APN consistently performs excellently. The outstanding performance of APN can be attributed to its innovative use of adaptive patching, which generates high-quality and regular initial patch representations. Building on these representations, APN employs a streamlined query module to effectively integrate contextual information, ensuring a lightweight and efficient framework design.
@@ 4.3 Ablation Studies
To validate the contributions of each key component of our APN model, we conduct ablation experiments—see Table 2. We have the following observations: 1) w/o Adaptive Patching: Replacing the learnable boundaries in TAPA with fixed time windows leads to performance degradation across all datasets, demonstrating the importance of the adaptive boundary learning mechanism. 2) w/o Weighted Aggregation: Replacing the time membership-based weighted averaging in TAPA with a simple arithmetic average of observations within the boundaries results in a significant performance drop. This highlights the core value of the time membership weighting strategy in synergizing with dynamic boundaries to further refine information and balance the information density within patches. 3) w/o Query-Based Patch Aggregation: Replacing the query-based intra-variable patch aggregation module with a simple linear mapping also leads to a notable performance loss, indicating that even with high-quality initial patch representations, effective context information integration remains crucial.
@@
Dataset PhysioNet MIMIC Human Activity USHCN
Metric MSE×$\times 10^{-3}$ MAE×$\times 10^{-2}$ MSE×$\times 10^{-2}$ MAE×$\times 10^{-2}$ MSE×$\times 10^{-3}$ MAE×$\times 10^{-2}$ MSE×$\times 10^{-1}$ MAE×$\times 10^{-1}$
w/o Adaptive Patching 4.58 $\pm$ 0.05 3.45 $\pm$ 0.03 1.65 $\pm$ 0.02 6.94 $\pm$ 0.04 2.78 $\pm$ 0.01 3.11 $\pm$ 0.00 4.38 $\pm$ 0.05 3.07 $\pm$ 0.04
w/o Weighted Aggregation 9.92 $\pm$ 0.11 5.81 $\pm$ 0.07 1.64 $\pm$ 0.03 6.96 $\pm$ 0.05 4.36 $\pm$ 0.12 4.28 $\pm$ 0.03 4.19 $\pm$ 0.02 2.95 $\pm$ 0.03
w/o Query-based Patch Aggregation 9.92 $\pm$ 0.10 4.71 $\pm$ 0.09 3.57 $\pm$ 0.03 11.58 $\pm$ 0.10 2.81 $\pm$ 0.02 3.18 $\pm$ 0.05 4.51 $\pm$ 0.04 3.52 $\pm$ 0.02
APN (ours) 4.48 $\pm$ 0.06 3.43 $\pm$ 0.02 1.61 $\pm$ 0.02 6.82 $\pm$ 0.05 2.77 $\pm$ 0.00 3.10 $\pm$ 0.01 4.09 $\pm$ 0.06 2.92 $\pm$ 0.05
@@ 4.4 Parameter Sensitivity
To gain a deeper understanding of the impact of key hyperparameters on the performance of the APN model, we conduct parameter sensitivity analyses, focusing primarily on the number of patches ($P$), the model’s hidden dimension ($D$), and the time encoding dimension ($D_{te}$)—see Table 3. We have the following observations: 1) From Figures 3(a) and 3(b), we observe that for most datasets, a larger $P$ does not necessarily lead to better performance. This suggests that an appropriate $P$ value can better balance the capture of local information and model complexity. 2) Figure 3(c) reveals the impact of the model’s hidden dimension ($D$). In general, a smaller hidden dimension may already be sufficient to capture effective information, while an excessively large dimension can sometimes lead to a slight performance decline. This indicates that the choice of $D$ should be comprehensively considered based on the scale and sparsity of the dataset. 3) Figure 3(d) shows the influence of the time encoding dimension ($D_{te}$). The results indicate that a moderate $D_{te}$ value can provide the model with effective temporal information. An excessively high time encoding dimension may introduce too much noise, interfering with the learning of core features. 4) The APN model exhibits a certain sensitivity to the selection of these core hyperparameters, but it generally achieves competitive results within a reasonable range. In practical applications, appropriate tuning should be performed based on the specific characteristics of the dataset.
@@ 4.5 Scalability and Efficiency Analysis
4.6 Varying Horizon
Table 3: Performance of varying observation and forecast horizons.
Algorithm History=3h, Forecast=45h History=12h, Forecast=36h History=36h, Forecast=12h History=45h, Forecast=3h
MSE×$\times 10^{-3}$ MAE×$\times 10^{-2}$ MSE×$\times 10^{-3}$ MAE×$\times 10^{-2}$ MSE×$\times 10^{-3}$ MAE×$\times 10^{-2}$ MSE×$\times 10^{-3}$ MAE×$\times 10^{-2}$
DLinear 51.82 17.13 43.56 15.73 41.63 15.48 41.23 15.51
TimesNet 57.30 10.70 24.95 7.62 13.57 5.50 13.86 5.65
PatchTST 42.18 13.67 18.56 7.80 9.85 5.11 8.53 4.64
Crossformer 9.48 5.86 8.57 5.70 5.70 4.47 5.33 4.44
Graph Wavenet 9.43 5.86 7.23 4.82 4.71 3.90 4.10 3.73
MTGNN 9.83 5.95 7.48 5.01 5.08 3.99 5.22 4.19
StemGNN 8.70 5.37 7.46 4.84 6.65 4.69 5.47 4.56
CrossGNN 10.44 6.56 7.97 5.37 6.87 4.73 4.80 4.24
FourierGNN 9.59 5.61 7.95 4.99 6.35 4.61 5.37 4.34
GRU-D 8.18 4.99 6.89 4.55 4.42 3.66 4.44 3.79
SeFT 9.78 5.55 9.30 5.41 9.15 5.15 8.76 5.57
RainDrop 10.47 5.72 9.89 5.62 9.70 5.40 9.28 5.62
Warpformer 8.48 5.13 7.57 4.83 5.60 4.09 6.44 4.67
mTAND 8.45 5.23 7.11 4.67 5.71 4.17 5.44 4.33
Latent-ODE 8.25 5.04 7.20 4.69 6.70 4.36 7.10 5.33
CRU 9.20 5.38 9.20 5.31 9.50 5.41 11.60 6.98
Neural Flow 8.30 4.99 8.50 5.27 7.70 4.68 7.40 5.10
T-PATCHGNN 8.01 4.87 6.48 4.19 4.14 3.31 3.69 3.25
TimeCHEAT 9.43 5.41 6.64 4.57 5.26 4.04 2.79 3.35
APN (ours) 8.79 4.90 5.97 3.94 3.91 3.08 2.51 2.65
To further evaluate the robustness of APN under varying combinations of observation history lengths and forecasting horizons, we conduct experiments on the PhysioNet dataset—as shown in Table 3. The results demonstrate that APN exhibits strong predictive capabilities and broad adaptability across multiple "History-Forecast" settings. In scenarios where historical information is relatively sufficient or the prediction span is moderate, APN achieves optimal performance in both MSE and MAE metrics, leading comprehensively. Under extreme settings with extremely limited historical information and long prediction spans, APN’s performance is comparable to several strong baseline models, ranking second only to T-PATCHGNN. Considering the significant challenges faced by all models under such extreme sparsity conditions, APN’s TAPA mechanism, which relies on learning adaptive features from data, may not fully realize its potential when historical information is scarce. Nevertheless, APN still demonstrates competitive robustness.
@@ 5 Conclusion
This paper addresses the challenges of IMTS forecasting by proposing a general and efficient APN. The core of APN lies in its novel TAPA module. Through learned dynamically adjustable patch boundaries and a time-aware weighting strategy, TAPA adaptively transforms irregular sequences into a fixed number of high-quality, regularized patch representations with highly condensed information, operating in a channel-independent manner. Leveraging the superior initial representations generated by TAPA, APN employs an exceptionally lean query module to integrate historical information and completes predictions via a shallow MLP. Extensive experiments on multiple public IMTS benchmark datasets demonstrate that APN significantly outperforms existing state-of-the-art methods in both prediction accuracy and computational efficiency.
Appendix A Detailed Experimental Setup
A.1 Dataset Details
To comprehensively and rigorously evaluate the predictive performance of our proposed APN model, we conduct extensive experiments on four real-world Irregular Multivariate Time Series (IMTS) datasets. These datasets are selected for their widespread use in benchmarking and public availability, representing typical characteristics of IMTS across different domains. Table 4 summarizes the key statistical features of these datasets.
For all datasets, samples are randomly partitioned into training, validation, and test sets, adhering to the standard 60%:20%:20% ratio. Regarding data preprocessing schemes (e.g., missing value handling, normalization methods, timestamp conversion) and prediction task configurations (e.g., historical observation window length, future prediction horizon length), we strictly follow established mainstream experimental protocols, particularly referencing benchmark settings established in related works such as T-PatchGNN [Zhang et al., 2024], to ensure the fairness and comparability of our experimental results. Next, we detail the characteristics of each dataset and its specific application in this study.
Table 4: Dataset statistics.
Dataset # Variable Avg # Obs. Avg Length Max Length
PhysioNet 41 10.7 75 216
MIMIC 96 1.5 46.1 643
Human Activity 12 30.2 120.6 130
USHCN 5 36.1 163.9 214
  • PhysioNet [Ansari et al., 2023]: Physionet [^1] https://archive.physionet.org/challenge/2012/ dataset contains time series data collected from 12,000 ICU patients. Each record details 41 clinical physiological indicators during the first 48 hours after ICU admission, with irregular sampling intervals and a significant amount of missing values. Following the standard settings of [Ansari et al., 2023, Rubanova et al., 2019], we use the observation data from the first 24 hours after each patient’s admission to predict their physiological indicator values for the subsequent 24 hours. Original data and access methods are available on the official PhysioNet website.
  • MIMIC [Naiman et al., 2024]: MIMIC [^2] https://physionet.org/content/mimiciii/1.4/ dataset is a large, publicly available database of de-identified critical care medical information. Following the commonly used preprocessing procedures in studies such as [Bilos et al., 2021, Naiman et al., 2024], we filter and process records to obtain 23,457 patient entries. Each record contains 96 irregularly sampled clinical variables from the first 48 hours after patient admission, also exhibiting significant irregularity and missingness. The experimental task is set to utilize data from the first 24 hours to predict variable values within the subsequent 24 hours. Access to this dataset requires certification and adherence to data use agreements via PhysioNet.
  • Human Activity [Zhang et al., 2023]: Human Activity [^3] https://archive.ics.uci.edu/dataset/196/localization+data+for+person+activity dataset comprises 12 motion-related variables (e.g., 3D acceleration, 3D gyroscope) recorded by sensors worn on different body parts (wrist, chest, ankle) of test subjects, with irregular sampling times. To construct IMTS samples suitable for forecasting tasks, we segment the original long sequences into clips of 4000 milliseconds and use the observation data from the first 3000 milliseconds to predict sensor readings for the next 1000 milliseconds. Through this process, we obtain a total of 5,400 IMTS samples to evaluate the model’s capability in biomechanical signal prediction.
  • USHCN [Ansari et al., 2023]: The USHCN dataset [^4] https://www.osti.gov/biblio/1394920 contains daily data for 5 core climate variables (daily maximum temperature, daily minimum temperature, daily precipitation, daily snow depth, daily snowfall) recorded from 1,114 weather stations across the United States between 1996 and 2000. This dataset is characterized by its long time span and naturally occurring spatial and temporal irregularities and data missingness due to station maintenance, equipment failure, etc. Following the experimental setup of [Ansari et al., 2023, Rubanova et al., 2019], we process the data for the task of using climate records from the past 24 months to predict climate conditions for the next 1 month, generating a total of 26,736 IMTS samples to test the model’s predictive performance on long-period, sparse climate data.
A.2 Baseline Model Details
To construct a comprehensive and convincing benchmark for the IMTS forecasting task, we integrate nineteen relevant baseline models for meticulous comparison. In selecting and evaluating these baseline models, we adhere to established mainstream experimental protocols [Zhang et al., 2024]. Specifically, we adapt state-of-the-art models from regular multivariate time series (RMTS) forecasting, IMTS classification, and IMTS imputation for the current IMTS forecasting task by appropriate modifications (e.g., altering the output layer for prediction tasks and incorporating timestamps and mask information as input features).
To ensure fairness and reliability of comparison, for some baseline models that have been rigorously evaluated in recent related works such as T-PatchGNN [Zhang et al., 2024], we prioritize referencing and reporting their publicly available optimal experimental results. For other models, or when evaluation under specific experimental settings is required, we reproduce them based on their official open-source implementations within a unified and consistent experimental framework. During reproduction, we conduct exploratory searches around the hyperparameter configurations recommended in the original papers and select the optimal parameter combinations based on performance on the validation set to report their best performance. This approach, combining references to existing SOTA results with careful reproduction, aims to provide a broad yet rigorous performance reference system for our APN model.
Next, we detail the core ideas, key mechanisms, and reference code sources for each baseline model:
Regular MTSF models:
  • DLinear [Zeng et al., 2023]: Models the trend and periodicity/residual terms of a time series using two independent single-layer linear networks, respectively, and then sums them to obtain the final prediction. Its design philosophy is that simple and effective linear models remain competitive in many forecasting tasks. Code Reference: https://github.com/cure-lab/LTSF-Linear
  • TimesNet [Wu et al., 2023]: Innovatively transforms one-dimensional time series into two-dimensional tensors to capture multi-periodicity. It identifies major periods via Fourier transform and reshapes sub-sequences of different periods into 2D tensors (with period-based "columns" and intra-period "rows"), then uses a 2D convolutional network to extract features, aiming to uniformly handle series with complex periodicities. Code Reference: https://github.com/thuml/TimesNet
  • PatchTST [Nie et al., 2023]: Segments each univariate time series into fixed-length non-overlapping patches, which serve as input tokens for a Transformer. It employs a Channel Independence strategy, modeling each variable independently, and learns temporal dependencies through a shared Transformer backbone. Code Reference: https://github.com/yuqinie/PatchTST
  • Crossformer [Wang et al., 2022]: A Transformer variant with a two-stage attention mechanism. The first stage captures intra-sequence temporal dependencies through dimension embedding and temporal self-attention; the second stage explicitly models interactions between different variables (dimensions) via an innovative Dimension-Segment-Wise Attention. Code Reference: https://github.com/Thinklab-AI/Crossformer
  • Graph Wavenet [Wu et al., 2019]: Combines Graph Convolutional Networks (GCN) with stacked dilated causal convolutions. It uses an adaptive adjacency matrix to learn dynamic spatial dependencies between variables and efficiently captures long-range temporal dependencies through dilated convolutions. Code Reference: https://github.com/nnzhan/Graph-WaveNet
  • MTGNN [Wu et al., 2020]: Proposes a general framework comprising a graph learning module, a graph convolution module, and a temporal convolution module. Its graph learning module adaptively infers unidirectional relationships between variables through node embedding and metric learning. Code Reference: https://github.com/nnzhan/MTGNN
  • StemGNN [Cao et al., 2020]: First transforms spatio-temporal graph data into the spectral domain (via Graph Fourier Transform and Discrete Fourier Transform), then designs modules in the spectral domain to capture spatial and temporal dependencies, and finally obtains predictions through inverse transformation. Code Reference: https://github.com/microsoft/StemGNN
  • CrossGNN [Huang et al., 2023]: To handle noise in multivariate time series, this model first generates multiple time series copies at different noise levels (data augmentation), then performs information interaction and feature refinement via cross-scale GNN and cross-variable GNN. Code Reference: https://github.com/hqh0728/CrossGNN
  • FourierGNN [Yi et al., 2023]: Treats multivariate time series from a pure graph perspective, where each time series is a node on the graph, and different time points are node features. Graph neural operations are performed in the Fourier domain to achieve prediction. Code Reference: https://github.com/aikunyi/FourierGNN
IMTS classification/imputation models:
  • GRU-D [Che et al., 2018]: An improved GRU model that models the influence of observation intervals through an introduced learnable time decay term and handles missing values via feature-level imputation to dynamically update hidden states. Code Reference: https://github.com/zhiyongc/GRU-D
  • SeFT (Set Functions for Time Series) [Horn et al., 2020]: Treats observation points of an irregular time series (typically as (time, value, variable ID) tuples) as an unordered set and models them using set functions like Deep Sets, naturally handling irregularity and variable length. Code Reference: https://github.com/mims-harvard/Raindrop
  • RainDrop [Zhang et al., 2022]: Models irregularly sampled data from a sensor network as a graph, where each observation point is a node. It captures spatio-temporal dependencies via graph attention networks and temporal attention mechanisms, generating interpolations or predictions for arbitrary query time points. Code Reference: https://github.com/mims-harvard/Raindrop
  • Warpformer [Zhang et al., 2023]: A multi-scale modeling approach for irregular clinical time series. Its core is the introduction of a learnable dynamic time warping module that adaptively aligns time series of different scales and sampling rates to a reference timeline before feeding them into a Transformer for representation learning. Code Reference: https://github.com/imJiawen/Warpformer
  • mTAND [Shukla and Marlin, 2021]: A multi-time attention network designed for irregularly sampled time series. It learns embedding representations for continuous timestamps and generates fixed-length context representations for variable-length input sequences, used for interpolation or classification. Code Reference: https://github.com/reml-lab/mTAN
IMTSF models:
  • Latent ODEs [Rubanova et al., 2019]: Compresses historical observation sequences (potentially variable-length and irregular) into a latent initial state via an RNN encoder, then evolves this state in latent space using a Neural ODE (an ordinary differential equation parameterized by a neural network), and finally generates predictions through a decoder. Code Reference: https://github.com/YuliaRubanova/latent-ode
  • CRU (Continuous Recurrent Units) [Schirmer et al., 2022]: Proposes a continuous-time recurrent unit whose hidden state evolution is driven by a linear stochastic differential equation, with irregularly spaced observation updates handled via Kalman filtering theory. Code Reference: https://github.com/boschresearch/Continuous-Recurrent-Units
  • Neural Flows [Bilos et al., 2021]: Reformulates the solving process of Neural ODEs as a continuous normalizing flow, aiming to improve numerical solution efficiency and stability by directly modeling the solution function of the ODE rather than its derivative. Code Reference: https://github.com/mbilos/neural-flows-experiments
  • T-PATCHGNN [Zhang et al., 2024]: Specifically designed for IMTS forecasting. Its core is to transform irregular univariate sequences into time-aligned patch sequences via a transformable patching method, then combine Transformers to capture intra-patch temporal dependencies and utilize time-adaptive graph neural networks to model dynamic inter-variable dependencies. Code Reference: https://github.com/usail-hkust/t-PatchGNN
  • TimeCHEAT [Liu et al., 2025]: Proposes a channel harmony strategy, employing channel-dependent graph convolutions within patches and channel-independent Transformers between patches to balance information processing across different channels. Code Reference: https://github.com/Alrash/TimeCHEAT
For all baseline models, we endeavor to use the official implementations provided by their authors or widely recognized community implementations. Regarding hyperparameter settings, we first refer to the recommendations in the original papers and then adjust them based on performance on the validation set to ensure fair comparison.
Appendix B Limitation
Although the APN method demonstrates its effectiveness, several aspects warrant further exploration:
  • Limited Scope of Validation Datasets: APN’s current validation primarily focuses on specific types of real-world datasets. Its adaptability and performance on more diverse domains, such as high-frequency financial data, industrial data with strong periodicities, or traffic flow with complex covariates, require further investigation to fully assess its generalization capabilities.
  • Performance in Extreme Scenarios: In situations with exceptionally short observation histories and very long prediction horizons, APN’s performance can be suboptimal. This is potentially because the data-driven adaptive mechanism struggles to converge effectively with extremely scarce information. Future work could explore incorporating prior knowledge or regularization to enhance robustness under such data-sparse conditions.

### v2

@@ 3.4 Query-based Weighted Aggregation
The TAPA module delivers a sequence of high-quality, regularized patch representations $H^{n}=[h^{n}_{1},\dots,h^{n}_{P}]$. This successfully concludes the first stage of our pipeline: handling irregularity. The second stage, contextualization and forecasting, can now proceed with remarkable efficiency. To this end, we employ a concise query-based aggregation mechanism to distill the entire historical sequence into a single, potent context vector $H_{c}^{n}$.
First, to preserve the sequential order of the patches, standard positional encodings ($PE$) are added, yielding position-aware representations $H_{pe}^{n}=H^{n}+PE$. Then, instead of resorting to complex inter-patch interactions like multi-head self-attention, we introduce a single learnable query vector $q^{n}\in\mathbb{R}^{D}$ for each channel. This query acts as a task-specific lens, dynamically assessing the importance of each patch for the final forecast. The importance scores $s_{p}^{n}$ are computed via a simple dot product, normalized into weights $\beta^{n}$ via Softmax, and used to compute the final context vector $H_{c}^{n}$ as a weighted sum:
@@
The resulting summary representations for all channels, $[H_{c}^{1},\dots,H_{c}^{N}]$, form the final representation matrix $H_{final}\in\mathbb{R}^{N\times D}$. A layer normalization is then applied to stabilize the input for the decoder. This lightweight aggregation mechanism is a direct testament to our design philosophy: by front-loading the complexity into TAPA, the subsequent modules can be elegantly simple yet powerful.
@@ 3.5 Forecasting Decoder
The final step of the APN framework is the ultimate demonstration of its architectural elegance. Having distilled the entire irregular history into a powerful, fixed-size representation $H_{final}^{n}$, the forecasting task becomes remarkably straightforward.
The prediction for a query time $\tau_{k}$ is made by a simple two-layer MLP decoder. This decoder takes the concatenated summary representation $H_{final}^{n}$ and the learnable temporal encoding of the query time, $TE(\tau_{k})$, as input to produce the final value $\hat{v}^{n}_{k}$:
@@
The ability to use such a simple decoder is not a limitation but a feature, underscoring the richness of the representation crafted by TAPA and the query aggregator.
The model is trained end-to-end by minimizing the standard Mean Squared Error (MSE) loss between the predicted values $\hat{v}^{n}_{k}$ and the ground truth values $v^{n}_{k}$ across all channels and query points:
$\mathcal{L}=\frac{1}{N\cdot Q}\sum_{n=1}^{N}\sum_{k=1}^{Q}(\hat{v}^{n}_{k}-v^{n}_{k})^{2},$ (10)
where $Q$ is the number of query points for each channel.
@@ 4.1 Setup
Implementation Details: We implement APN and all related experiments using the PyTorch framework on a server equipped with an Intel(R) Xeon(R) Gold 6326 CPU (@ 2.90GHz) and an NVIDIA GeForce RTX 3090 GPU. For our APN model, we set the hidden dimension $D$ according to different datasets. Specifically, the hidden dimension $D$ is set to 64 for the PhysioNet and USHCN, while it is set to 16 for the Human Activity and MIMIC. The time encoding dimension $D_{te}$ is uniformly set to 10. The number of patches $P$ is adjusted based on the characteristics of the datasets: 20 for PhysioNet, 16 for MIMIC, 75 for Human Activity, and 4 for USHCN. The training batch size is set to 256 for all datasets. The model is trained using the Adam optimizer with an initial learning rate of $1\times 10^{-2}$. We train the model for a maximum of 200 epochs and employ an early stopping strategy with a patience of 50. To ensure the robustness of the results, we repeat each experiment 5 times using five different random seeds. We then report the mean and standard deviation of the performance metrics across these runs.
@@
[…]
@@ 4.2 Main results
We compare the APN model with nineteen baseline models on four challenging datasets—see Table 1. We have the following observations: 1) APN achieves leading prediction accuracy across the board. On all datasets, APN achieves optimal or highly competitive forecasting performance. Compared to the second-best performing model, T-PATCHGNN, APN achieves significant reductions in MSE and MAE metrics by approximately 9.63% and 5.24%, respectively. 2) APN demonstrates exceptional cross-domain generalization capability and robustness. Whether on IMTS datasets with different characteristics in healthcare (PhysioNet, MIMIC), biomechanics (Human Activity), or climate science (USHCN), APN consistently performs excellently. The outstanding performance of APN can be attributed to its TAPA module, which generates superior patch representations by directly performing a soft-aggregation over raw observations within adaptively learned temporal windows. This approach effectively captures salient local dynamics without the potential distortion from interpolation. Building on these high-quality representations, the streamlined query module and MLP decoder ensure a lightweight yet powerful framework.
@@ 4.3 Ablation Studies
We perform ablation studies to isolate and validate the contribution of each primary component of APN. The complete results are provided in the Appendix B.2, with a representative summary presented in Table 2. The experiments yield three key observations. First, removing the Adaptive Patching mechanism and reverting to fixed-size windows results in a consistent decline in performance, which highlights the value of dynamically learned boundaries. Second, ablating the Weighted Aggregation scheme for a simple average over hard-cut patches leads to the most significant performance degradation. This empirically confirms that our soft-aggregation design is critical for preventing information loss at patch boundaries (i.e., the coverage issue). Third, replacing the Query-Based Patch Aggregation with a basic linear layer also diminishes performance, proving that our sophisticated aggregation method is essential for integrating contextual information within variables.
Dataset PhysioNet MIMIC USHCN
[…]
w/o Adaptive
Patching
6.98 $\pm$ 0.05 2.05 $\pm$ 0.02 5.59 $\pm$ 0.05
w/o Weighted
Aggregation
9.92 $\pm$ 0.11 2.27 $\pm$ 0.03 5.38 $\pm$ 0.02
w/o Query-based
Patch Aggregation
9.92 $\pm$ 0.10 3.57 $\pm$ 0.03 5.21 $\pm$ 0.04
APN (ours) 4.48 $\pm$ 0.06 1.61 $\pm$ 0.02 4.52 $\pm$ 0.06
@@ 4.4 Parameter Sensitivity
To gain a deeper understanding of the impact of key hyperparameters on the performance of the APN model, we conduct parameter sensitivity analyses, focusing primarily on the number of patches ($P$), the model’s hidden dimension ($D$), and the time encoding dimension ($D_{te}$)—see Figure 3. We have the following observations: 1) From Figures 3(a) and 3(b), we observe that for most datasets, a larger $P$ does not necessarily lead to better performance. This suggests that an appropriate $P$ value can better balance the capture of local information and model complexity. 2) Figure 3(c) reveals the impact of the model’s hidden dimension ($D$). In general, a smaller hidden dimension may already be sufficient to capture effective information, while an excessively large dimension can sometimes lead to a slight performance decline. This indicates that the choice of $D$ should be comprehensively considered based on the scale and sparsity of the dataset. 3) Figure 3(d) shows the influence of the time encoding dimension ($D_{te}$). The results indicate that a moderate $D_{te}$ value can provide the model with effective temporal information. An excessively high time encoding dimension may introduce too much noise, interfering with the learning of core features. 4) The APN model exhibits a certain sensitivity to the selection of these core hyperparameters, but it generally achieves competitive results within a reasonable range. In practical applications, appropriate tuning should be performed based on the specific characteristics of the dataset.
@@ 5 Conclusion
This paper addresses the challenges of IMTS forecasting by proposing a general and efficient framework, APN. The core of APN lies in its novel Time-Aware Patch Aggregation (TAPA) module. This module introduces an aggregation-based paradigm, distinct from methods that rely on fixed patching or interpolation. By learning dynamic boundaries and applying a time-aware soft-weighting strategy, TAPA directly aggregates information from all raw observations to generate high-quality, regularized patch representations. This design choice not only ensures full data coverage and robustly handles information density variations, but also enables a streamlined and efficient overall architecture. Leveraging these superior representations, APN employs a concise query module and a shallow MLP to make predictions. Extensive experiments on multiple public IMTS benchmark datasets demonstrate that APN significantly outperforms existing state-of-the-art methods in both prediction accuracy and computational efficiency.